# Lagrangian chaos and unique ergodicity for stochastic primitive equations

### v1

@@ 6.4.1. Construction of the control
$\gamma^{(w)}_{(\mathbf{k},\ell)}\stackrel{{\scriptstyle{\rm def}}}{{=}}\left\{\begin{aligned} -&\,k_{z}^{-1}(k_{x},k_{y})\cdot\gamma_{(\mathbf{k},\ell)}&\ &\text{ if }\mathbf{k}=(k_{x},k_{y},k_{z})\text{ satisfies }k_{z}\neq 0,\\ &0&\ &\text{ otherwise. }\end{aligned}\right.$
Letting $\gamma^{(u)}_{(\mathbf{k},\ell)}\stackrel{{\scriptstyle{\rm def}}}{{=}}(\gamma_{(\mathbf{k},\ell)},\gamma^{(w)}_{(\mathbf{k},\ell)})\in{\mathbb{R}}^{3}$, we obtain that the velocity field $u$ corresponding to an horizontal velocity $v\in\mathcal{H}_{\mathrm{L}}$ decomposed as (6.57) has the form
@@
$u(\mathbf{x})=\sum_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}v_{(\mathbf{k},\ell)}\gamma^{(u)}_{(\mathbf{k},\ell)}e_{\mathbf{k}}\ \ \text{ for }\mathbf{x}\in\mathbb{T}^{3}.$
Therefore, for each $\xi\in\mathbb{S}^{2}$,
@@
$\Pi_{\xi}\big{[}\nabla u(\mathbf{x})\xi\big{]}=\sum_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}v_{(\mathbf{k},\ell)}(\mathbf{k}\cdot\xi)e_{-\mathbf{k}}(\mathbf{x})\,\Pi_{\xi}\gamma_{(\mathbf{k},\ell)}^{(u)}\ \ \text{ for }\mathbf{x}\in\mathbb{T}^{3}.$
Now, we return to the proof of (6.56). With the above notation, for $j=(\mathbf{k},\ell)\in\mathcal{J}_{\mathrm{L}}$, by (6.50) and noticing that
@@
it follows that
@@
(6.58) $D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j}=q_{(\mathbf{k},\ell)}\begin{pmatrix}0\\ e_{\mathbf{k}}(\mathbf{x})\gamma^{(u)}_{(\mathbf{k},\ell)}\\ (\mathbf{k}\cdot\xi)e_{-\mathbf{k}}(\mathbf{x})\Pi_{\xi}\gamma_{(\mathbf{k},\ell)}^{(u)}\\ 0\end{pmatrix}.$
Note that
$\mathrm{span}\big{\{}\gamma_{(\mathbf{k},1)}^{(u)},\gamma_{(\mathbf{k},2)}^{(u)}\big{\}}=\big{\{}\mathbf{k}^{\prime}\in{\mathbb{R}}^{3}\,:\,\mathbf{k}^{\prime}\cdot\mathbf{k}=0\big{\}}$
for all $\mathbf{k}=(k_{x},k_{y},k_{z})$ such that $k_{z}\neq 0$ (the latter condition is needed since the vector $(0,0,1)$ that serves as $\gamma_{\mathbf{k},2}^{(u)}$ in case $k_{z}=0$ is not included in $\mathcal{Z}$, see Subsection 2.2.2). Thus, by the argument in [16, Lemma 5.3], for all for all $\mathbf{x}\in\mathbb{T}^{3}$ and $\xi\in\mathbb{S}^{2}$, it holds that
$\mathrm{span}\big{\{}e_{\mathbf{k}}(\mathbf{x})\gamma^{(u)}_{(\mathbf{k},\ell)}\,:\,(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}\big{\}}$ $={\mathbb{R}}^{3},$
$\mathrm{span}\big{\{}(\mathbf{k}\cdot\xi)e_{-\mathbf{k}}(\mathbf{x})\Pi_{\xi}\gamma_{(\mathbf{k},\ell)}^{(u)}\,:\,(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}\big{\}}$ $=T_{\xi}\mathbb{S}^{2},$
and the estimate (6.56) holds with a constant (possibly) depending on $(\mathbf{x},\xi)$. By continuity and compactness of $\mathbb{T}^{3}\times\mathbb{S}^{2}$, the lower bound (6.56) can be made uniform in $(\mathbf{x},\xi)$. Finally, let us remark that the independence of the constant $c_{0}$ in (6.56) on $v\in\mathbb{H}^{\sigma}(\mathbb{T}^{3})$ is due to the $v$-independence of the RHS(6.58). ∎
Before going further, let us point out the following observation.
Remark 6.12 (Hörmander condition for the low-mode SDEs).
As in [16, Section 5], the result of Lemma 6.10 yields that the system
(6.59) $\partial_{t}\mathbf{x}_{t}$ $=u_{t}(\mathbf{x}_{t}),$
(6.60) $\partial_{t}\xi_{t}$ $=\Pi_{\xi_{t}}\big{[}\nabla u_{t}(\mathbf{x}_{t})\xi_{t}\big{]},$
(6.61) $\partial_{t}v_{t}^{(\mathbf{k},\ell)}$ $=-|\mathbf{k}|^{2}v_{t}^{(\mathbf{k},\ell)}+q_{(\mathbf{k},\ell)}\dot{W}_{t}^{(\mathbf{k},\ell)}\ \ \ \text{ for }\ (\mathbf{k},\ell)\in\mathcal{Z}_{0},$
where $u_{t}=(v_{t},w(v_{t}))$, satisfies the Hörmander condition. The reader is referred to [16, Section 5] for further details and references.
Finally, we prove Lemma 6.10.
@@
In particular, as in Case 1 of [16, Lemma 6.15], we can use the regularizing $\eta$ process to span the missing directions $(\mathbf{x},\xi)$. More precisely, for $j\in\{1,\dots,6\}$, since $D_{\mathrm{L}}F_{\mathrm{L}}(V)\mathbf{q}_{j}=\frac{\mathrm{d}}{\mathrm{d}\varepsilon}H((\xi,\eta)+\varepsilon a_{j})|_{\varepsilon=0}$, it holds that
$|\langle D_{\mathrm{L}}F_{\mathrm{L}}(V)\mathbf{q}_{j},h\rangle|\geq\frac{1}{(1+|\eta|^{2})^{5/2}}|\langle a_{j},h\rangle|\geq\frac{1}{(1+|\eta|)^{5}}|\langle a_{j},h\rangle|.$
@@
Hence, in case $j=(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}$, arguing as for (6.58), we have
@@
$-\frac{1}{r}\,\chi^{\prime}(\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}/{r})\frac{q_{(\mathbf{k},\ell)}v_{(\mathbf{k},\ell)}}{\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}H(\xi,\eta).$
where $v_{(\mathbf{k},\ell)}=\int_{\mathbb{T}^{3}}v(\mathbf{x})\cdot\gamma_{(\mathbf{k},\ell)}e_{\mathbf{k}}(\mathbf{x})\,\mathrm{d}\mathbf{x}$. Now, we apply the operator $\langle\cdot,h\rangle$, and we estimate each term separately. Firstly, by (6.53), $\|u\|_{\mathbb{H}^{\sigma}}\leq 2r$ and Lemma 5.2,
$|\langle D_{\mathrm{L}}F^{{\rm vel}}_{\mathrm{L}}(v)\mathbf{q}_{j},h\rangle|=|\langle D_{\mathrm{L}}F^{{\rm vel}}_{\mathrm{L}}(v)\mathbf{q}_{j},v_{\mathrm{L}}^{(h)}\rangle|\lesssim_{r}|v_{\mathrm{L}}^{(h)}|\lesssim\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle\mathbf{q}_{j},h\rangle|.$
@@
$\frac{1}{r}\Big{|}\chi^{\prime}\Big{(}\frac{\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}{r}\Big{)}\frac{q_{(\mathbf{k},\ell)}v_{(\mathbf{k},\ell)}}{\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}H(\xi,\eta)\Big{|}\leq\frac{c_{0}}{2}|h|.$
@@
$|\langle[A_{\mathrm{L}}-D_{\mathrm{L}}F_{\mathrm{L}}(V)]\mathbf{q}_{j},h\rangle|\geq|\langle D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j},h\rangle|-K\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle\mathbf{q}_{j},h\rangle|-\frac{c_{0}}{2}|h|,$
and therefore
$\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle[A_{\mathrm{L}}-D_{\mathrm{L}}F_{\mathrm{L}}(V)]\mathbf{q}_{j},h\rangle|\geq\frac{c_{0}}{2}|h|-K\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle\mathbf{q}_{j},h\rangle|.$
The claim now follows from (6.56) in Lemma 6.11. ∎

### v2

@@ 6.4.1. Construction of the control
$\gamma^{(w)}_{(\mathbf{k},\ell)}$ =$\stackrel{{\scriptstyle{\rm def}}}{{=}}-\,k_{z}^{-1}(k_{x},k_{y})\cdot a_{\ell}\ \ \text{ if }k_{z}\neq 0,\quad\text{ or }\gamma^{(w)}_{(\mathbf{k},\ell)}\stackrel{{\scriptstyle{\rm def}}}{{=}}0\ \text{ otherwise. }$.
Note that $\gamma^{(w)}_{(\mathbf{k},\ell)}=\gamma^{(w)}_{(-\mathbf{k},\ell)}$ by symmetry of $\gamma_{(\cdot,\ell)}$. Letting
@@
$\gamma^{(u)}_{(\mathbf{k},\ell)}\stackrel{{\scriptstyle{\rm def}}}{{=}}(\gamma_{(\mathbf{k},\ell)},\gamma^{(w)}_{(\mathbf{k},\ell)})\in{\mathbb{R}}^{3}\quad\text{ and }\quad\beta^{(u)}_{(\mathbf{k},\ell)}\stackrel{{\scriptstyle{\rm def}}}{{=}}(0,0,\gamma^{(w)}_{(\mathbf{k},\ell)})\in{\mathbb{R}}^{3},$
we obtain that the velocity field $u$ corresponding to an horizontal velocity $v\in\Pi_{\mathrm{L}}(\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3}))$ decomposed as (6.57) has the form
$u(\mathbf{x})=\sum_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}v_{(\mathbf{k},\ell)}\big{(}\gamma^{(u)}_{(\mathbf{k},\ell)}e_{\mathbf{k}}(\mathbf{x})-\beta^{(u)}_{(\mathbf{k},\ell)}e_{\mathbf{k}}(\mathbf{x}_{0})\big{)}.$
Therefore, for each $\xi\in\mathbb{S}^{2}$ and $\mathbf{x}\in\mathcal{O}$,
@@
$\Pi_{\xi}\big{[}\nabla u(\mathbf{x})\xi\big{]}\textstyle=\sum_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}v_{(\mathbf{k},\ell)}\big{(}(\mathbf{k}\cdot\xi)e_{-\mathbf{k}}(\mathbf{x})\,\Pi_{\xi}\gamma_{(\mathbf{k},\ell)}^{(u)}-(\mathbf{k}\cdot\xi_{0})e_{-\mathbf{k}}(\mathbf{x}_{0})\,\Pi_{\xi}\beta_{(\mathbf{k},\ell)}^{(u)}\big{)}.$
where $\xi_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}(\xi_{x},\xi_{y},0)$. With the above preparation, we first establish Lemma 6.11, and then turn to Lemma 6.10.
Proof of Lemma 6.11.
The estimate (6.55) follows clearly from $\inf_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}q_{(\mathbf{k},\ell)}>0$ and (6.50). Next, we focus on (6.56). Since for all $j=(\mathbf{k},\ell)\in\mathcal{J}_{\mathrm{L}}$ and $\mathbf{x}\in\mathcal{O}$
@@
it follows that $D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j}$ is given by
(6.58) $q_{(\mathbf{k},\ell)}\begin{pmatrix}0\\ \gamma^{(u)}_{(\mathbf{k},\ell)}e_{\mathbf{k}}(\mathbf{x})-\beta^{(u)}_{(\mathbf{k},\ell)}e_{\mathbf{k}}(\mathbf{x}_{0})\\ \Pi_{\xi}\big{[}(\mathbf{k}\cdot\xi)e_{-\mathbf{k}}(\mathbf{x})\,\gamma_{(\mathbf{k},\ell)}^{(u)}-(\mathbf{k}\cdot\xi_{0})e_{-\mathbf{k}}(\mathbf{x}_{0})\,\beta_{(\mathbf{k},\ell)}^{(u)}\big{]}\\ 0\end{pmatrix}.$
For exposition convenience, we now split the proof into two steps.
Step 1: There exist positive constants $r_{1},c_{1}$ such that, for all $r\geq r_{1}$, $V=(v_{\mathrm{L}},\mathbf{x},\xi,\eta)\in\mathcal{H}_{\mathrm{L}}$ and $h=(h_{v},h_{\mathbf{x}},h_{\xi},h_{\eta})\in T_{V}\mathcal{H}_{\mathrm{L}}$,
(6.59) $\max_{j\in\mathcal{Z}_{\mathrm{L}}}|\langle D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j},h\rangle|\geq c_{1}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}|h_{\mathbf{x}}|.$
For $j=(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}$, we denote by $j^{\prime}$ the reflected index $j^{\prime}=(-\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}$. By standard computations, one can check that the vector $e_{\mathbf{k}}(\mathbf{x})D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j}+e_{-\mathbf{k}}(\mathbf{x})D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j^{\prime}}$ is given by
@@
$\Phi_{(\mathbf{k},\ell)}(z,\xi)\stackrel{{\scriptstyle{\rm def}}}{{=}}q_{(\mathbf{k},\ell)}\begin{pmatrix}0\\ \gamma^{(u)}_{(\mathbf{k},\ell)}-\beta^{(u)}_{(\mathbf{k},\ell)}\cos(2\pi k_{z}z)\\ \mathrm{sign}(\mathbf{k})\,[(\mathbf{k}\cdot\xi_{0})\,\Pi_{\xi}\beta_{(\mathbf{k},\ell)}^{(u)}\,\sin(2\pi k_{z}z)]\\ 0\end{pmatrix},$
where, as in (2.11), $\mathrm{sign}(\mathbf{k})=1$ if $\mathbf{k}\in{\mathbb{Z}}^{3}_{+}$ or $\mathrm{sign}(\mathbf{k})=0$ otherwise. In particular, the vector $\tfrac{1}{2}(\Phi_{(\mathbf{k},\ell)}(z,\xi)+\Phi_{(-\mathbf{k},\ell)}(z,\xi))$ has only one non-trivial component, namely the one in the $\mathbf{x}$-direction, which takes the form
$\phi_{(\mathbf{k},\ell)}(z)\stackrel{{\scriptstyle{\rm def}}}{{=}}\gamma^{(u)}_{(\mathbf{k},\ell)}-\beta^{(u)}_{(\mathbf{k},\ell)}\cos(2\pi k_{z}z).$
To prove (6.59), it is enough to show that, for all vectors $h\in{\mathbb{R}}^{3}$,
(6.60) $\max_{(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}}|\langle\phi_{(\mathbf{k},\ell)}(z),h\rangle|\gtrsim(|z|^{2}\wedge|1-z|^{2})|h|.$
To prove the above, let $\mathbf{k}_{1}=(1,0,1)$ and $\mathbf{k}_{2}=(1,-1,0)$. It is clear that
$\gamma_{(\mathbf{k}_{1},1)}=\begin{pmatrix}1\\ 0\\ \cos(2\pi z)-1\end{pmatrix},\qquad\gamma_{(\mathbf{k}_{1},2)}=\begin{pmatrix}0\\ 1\\ 0\end{pmatrix},\qquad\gamma_{(\mathbf{k}_{2},1)}=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ -1\\ 0\end{pmatrix}.$
Since $\mathrm{span}\{\gamma_{(\mathbf{k}_{1},2)},\gamma_{(\mathbf{k}_{2},1)}\}={\mathbb{R}}^{2}\times\{0\}$, the claim (6.60) follows by noticing that $(\mathbf{k}_{1},1),(\mathbf{k}_{1},2),(\mathbf{k}_{2},1)\in\mathcal{Z}_{\mathrm{L}}$ and $1-\cos(2\pi z)\gtrsim|z|^{2}\wedge|1-z|^{2}$ by Taylor’s expansion.
Step 2: There exist positive constants $r_{2},c_{2}$ such that, for all $r\geq r_{2}$, $V=(v_{\mathrm{L}},\mathbf{x},\xi,\eta)\in\mathcal{H}_{\mathrm{L}}$ and $h=(h_{v},h_{\mathbf{x}},h_{\xi},h_{\eta})\in T_{V}\mathcal{H}_{\mathrm{L}}$,
(6.61) $\max_{j\in\mathcal{Z}_{\mathrm{L}}}|\langle D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j},h\rangle|\geq c_{2}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}|h_{\xi}|.$
By (6.58) and the argument in Step 2, it is enough to show that
(6.62) $\mathrm{span}\big{\{}\Psi_{(\mathbf{k},\ell)}(z,\xi)\,:\,(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}\big{\}}=T_{\xi}\mathbb{S}^{2},$
with a corresponding lower bound, where
$\Psi_{(\mathbf{k},\ell)}(\mathbf{x},\xi)\stackrel{{\scriptstyle{\rm def}}}{{=}}\Pi_{\xi}\big{[}(\mathbf{k}\cdot\xi)e_{-\mathbf{k}}(\mathbf{x})\,\gamma_{(\mathbf{k},\ell)}^{(u)}-(\mathbf{k}\cdot\xi_{0})e_{-\mathbf{k}}(\mathbf{x}_{0})\,\beta_{(\mathbf{k},\ell)}^{(u)}\big{]}.$
Similar in Step 2, the vector $e_{-\mathbf{k}}(\mathbf{x})\Psi_{(\mathbf{k},\ell)}(\mathbf{k},\xi)-e_{\mathbf{k}}(\mathbf{x})\Psi_{(\mathbf{k},\ell)}(\mathbf{k},\xi)$ takes the form
$\psi_{(\mathbf{k},\ell)}(z,\xi)\stackrel{{\scriptstyle{\rm def}}}{{=}}\Pi_{\xi}\big{[}(\mathbf{k}\cdot\xi)\gamma_{(\mathbf{k},\ell)}^{(u)}-(\mathbf{k}\cdot\xi_{0})\,\beta_{(\mathbf{k},\ell)}^{(u)}\,\cos(2\pi k_{z}z)\big{]}.$
To prove (6.62), we distinguish two cases depending on a parameter $\delta>0$, which will be chosen depending on $\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})$. Below, we write $\xi=(\xi_{x},\xi_{y},\xi_{z})\in{\mathbb{R}}^{3}$, and for $y\in{\mathbb{R}}$, we let $\mathrm{sign}(y)=1$ if $y\geq 0$, or $\mathrm{sign}(y)=-1$ if $y<0$.
Claim 1: There exists $K>0$ such that (6.61) holds with RHS replaced by $c_{2}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{2}|h_{\xi}|$ provided $|\xi_{z}|\leq K(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{2}$. By a standard perturbation argument, in this case, it suffices to prove Claim 1 for $\xi_{z}=0$. Note that, as $\xi_{z}=0$ and $|\xi|=1$, either $\xi_{x}\geq\frac{1}{\sqrt{2}}$ or $\xi_{y}\geq\frac{1}{\sqrt{2}}$. In the following, we only discuss the case $\xi_{x}\geq\frac{1}{\sqrt{2}}$, as the other is similar. Let $\mathbf{k}_{1}=(\mathrm{sign}(\xi_{x}),\mathrm{sign}(\xi_{y}),0)$ and $\mathbf{k}_{2}=(\mathrm{sign}(\xi_{x}),0,1)$. Thus,
$\psi_{(\mathbf{k}_{1},1)}(z,\xi)=\sqrt{2}(|\xi_{x}|+|\xi_{y}|)\Pi_{\xi}\begin{pmatrix}\mathbf{k}_{1}^{\perp}\\ 0\end{pmatrix},\quad\psi_{(\mathbf{k}_{2},1)}(z,\xi)=|\xi_{x}|\,\Pi_{\xi}\begin{pmatrix}1\\ 0\\ -1+\cos(2\pi z)\end{pmatrix}.$
Due to $\xi_{z}=0$, it holds that $T_{\xi}{\mathbb{R}}^{2}=\mathrm{span}\{(\xi^{\perp},0)^{\top},(0,0,1)^{\top}\}$. Moreover, from $\mathbf{k}_{1}\cdot\xi\neq 0$ it follows that $\Pi_{\xi}(\mathbf{k}_{1}^{\perp},0)^{\top}=\mathrm{span}\{(\xi^{\perp},0)\}$. Now, in the case $\xi_{z}=0$, the claim (6.61) holds with the RHS replaced by $c_{2}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{2}|h_{\xi}|$. As noted earlier, by perturbation, this is sufficient to prove Claim 1.
Claim 2: Let $K>0$ be as in Claim 1. Then (6.61) holds for $|\xi_{z}|\geq K(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{2}$. Here we proceed as in Step 1 of the current proof. Consider
$\widetilde{\psi}_{(\mathbf{k},\ell)}(z,\xi)\stackrel{{\scriptstyle{\rm def}}}{{=}}(\mathbf{k}\cdot\xi)\gamma_{(\mathbf{k},\ell)}^{(u)}-(\mathbf{k}\cdot\xi_{0})\,\beta_{(\mathbf{k},\ell)}^{(u)}\,\cos(2\pi k_{z}z).$
Here, instead of (6.62), we prove $\mathrm{span}\big{\{}\widetilde{\psi}_{(\mathbf{k},\ell)}(z,\xi)\,:\,(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}\big{\}}={\mathbb{R}}^{3}$ with a corresponding lower bound as in (6.61). Indeed, for $\mathbf{k}_{1}=(\mathrm{sign}(\xi_{x}),0,\mathrm{sign}(\xi_{z}))$ and $\mathbf{k}_{2}=(0,\mathrm{sign}(\xi_{y}),\mathrm{sign}(\xi_{z}))$, one has
$\widetilde{\psi}_{(\mathbf{k}_{1},1)}(z,\xi)$ $=(|\xi_{x}|+|\xi_{z}|)\begin{pmatrix}1\\ 0\\ -\frac{\mathrm{sign}(\xi_{x})}{\mathrm{sign}(\xi_{z})}\end{pmatrix}+|\xi_{x}|\begin{pmatrix}1\\ 0\\ \frac{\mathrm{sign}(\xi_{x})}{\mathrm{sign}(\xi_{z})}\end{pmatrix}\cos(2\pi z),$
$\widetilde{\psi}_{(\mathbf{k}_{2},2)}(z,\xi)$ $=(|\xi_{y}|+|\xi_{z}|)\begin{pmatrix}0\\ 1\\ -\frac{\mathrm{sign}(\xi_{y})}{\mathrm{sign}(\xi_{z})}\end{pmatrix}+|\xi_{y}|\begin{pmatrix}0\\ 1\\ \frac{\mathrm{sign}(\xi_{y})}{\mathrm{sign}(\xi_{z})}\end{pmatrix}\cos(2\pi z),$
$\widetilde{\psi}_{(\mathbf{k}_{1},2)}(z,\xi)$ $=(|\xi_{x}|+|\xi_{z}|)\begin{pmatrix}0\\ 1\\ 0\end{pmatrix},\quad\text{ and }\quad\widetilde{\psi}_{(\mathbf{k}_{2},1)}(z,\xi)=(|\xi_{x}|+|\xi_{z}|)\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}.$
The above readily implies the estimate (6.61) as $|\xi_{z}|(1-\cos(2\pi z))\gtrsim(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}$.
The claim (6.61) of Step 2 hence follows from Claims 1 and 2. This finished the proof of Lemma 6.11. ∎
@@
In particular, as in Case 1 of [16, Lemma 6.15], we can use the regularising $\eta$ process to span the missing directions $(\mathbf{x},\xi)$. More precisely, for $j\in\{1,\dots,6\}$, since $D_{\mathrm{L}}F_{\mathrm{L}}(V)\mathbf{q}_{j}=\frac{\mathrm{d}}{\mathrm{d}\varepsilon}H(\mathbf{x},(\xi,\eta)+\varepsilon a_{j})|_{\varepsilon=0}$, it holds that
$|\langle D_{\mathrm{L}}F_{\mathrm{L}}(V)\mathbf{q}_{j},h\rangle|\geq\frac{(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}}{(1+|\eta|^{2})^{3/2}}|\langle a_{j},h\rangle|\geq\frac{(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}}{(1+|\eta|)^{3}}|\langle a_{j},h\rangle|.$
@@
Hence, in case $j=(\mathbf{k},\ell)\in\mathcal{Z}_{\mathrm{L}}$, we have
@@
$-\frac{1}{r}\,\chi^{\prime}\Big{(}\frac{\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}{r}\Big{)}\frac{q_{(\mathbf{k},\ell)}(v,\gamma_{(\mathbf{k},\ell)}e_{\mathbf{k}})_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}{\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}H(\mathbf{x},\xi,\eta).$
where $v_{(\mathbf{k},\ell)}=\int_{\mathbb{T}^{3}}v(\mathbf{x})\cdot\gamma_{(\mathbf{k},\ell)}e_{\mathbf{k}}(\mathbf{x})\,\mathrm{d}\mathbf{x}$. Now, we apply the operator $\langle\cdot,h\rangle$, and we estimate each term separately. As above, here $h=(h_{v},h_{\mathbf{x}},h_{\xi},h_{\eta})\in\mathbb{H}^{\sigma}_{0}\times{\mathbb{R}}^{3}\times T_{\xi}\mathbb{S}^{2}\times{\mathbb{R}}^{6}$. Firstly, by (6.53), $\|v\|_{\mathbb{H}^{\sigma}}\leq 2r$ and Lemma 5.2,
$|\langle D_{\mathrm{L}}F^{{\rm vel}}_{\mathrm{L}}(v)\mathbf{q}_{j},h\rangle|=|\langle D_{\mathrm{L}}F^{{\rm vel}}_{\mathrm{L}}(v)\mathbf{q}_{j},h_{v}\rangle|\lesssim_{r}|h_{v}|\lesssim\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle\mathbf{q}_{j},h\rangle|.$
@@
$\frac{1}{r}\sup_{j\in\mathcal{Z}_{\mathrm{L}}}\Big{|}\chi^{\prime}\Big{(}\frac{\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}{r}\Big{)}\frac{q_{(\mathbf{k},\ell)}(v,\gamma_{(\mathbf{k},\ell)}e_{\mathbf{k}})_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}{\|v\|_{\mathbb{H}^{\sigma}(\mathbb{T}^{3})}}H(\mathbf{x},\xi,\eta)\Big{|}\leq\frac{c_{0}}{2}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}.$
@@
|$\langle$| ⟨ $[A_{\mathrm{L}}-D_{\mathrm{L}}F_{\mathrm{L}}(V)]\mathbf{q}_{j},h\rangle|$
$\geq|\langle D_{\mathrm{L}}F^{{\rm pr}}_{\mathrm{L}}(v,\mathbf{x},\xi)\mathbf{q}_{j},h\rangle|-K\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle\mathbf{q}_{j},h\rangle|-\tfrac{1}{2}c_{0}\,(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}(|h_{\mathbf{x}}|+|h_{\xi}|),$
where $h_{\mathbf{x}}$ and $h_{\xi}$ are the $\mathbf{x}$- and $\xi$-components of $h\in T_{V}\mathcal{H}$, respectively. Additional, we used the fact that $H(\mathbf{x},\xi,\eta)$ has only non-trivial $\mathbf{x}$ and $\xi$-components.
Therefore, from (6.56), we get
$\max_{j\in\mathcal{Z}_{\mathrm{L}}}|\langle[A_{\mathrm{L}}-D_{\mathrm{L}}F_{\mathrm{L}}(V)]\mathbf{q}_{j},h\rangle|\geq\tfrac{1}{2}c_{0}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}(|h_{\mathbf{x}}|+|h_{\xi}|)-K\max_{j\in\mathcal{J}_{\mathrm{L}}}|\langle\mathbf{q}_{j},h\rangle|.$
The claim (6.52) now follows from (6.55) in Lemma 6.11. ∎